# Gradient-free neural topology optimization

### v1

@@ Appendix E Alternative configurations of the baseline VAE model
For the baseline VAE model we tested 4 different configurations: 2 different architecture configurations (see Table 2), and 2 different parameterizations of the latent space. The results are summarized in Figure 22.
@@
Figure 22: Cumulative probability vs relative error (w.r.t. the MMA solution) for different optimization methods, evaluated on smaller testing sets (10 problems each, unlike on the full results), repeated with 10 different random seeds. The distribution does not add up to 1 due to cut-off at 300%
@@
For optimization with VAE, we explore two different parameterizations of the latent space: 1) where the latent optimizer controls the whole input to the decoder ($\vec{z}$), and 2) where we utilize the information from the encoder, and the optimizer controls the additive term $\epsilon$, such that the input to the decoder $z=\mu+\sigma\epsilon$, where $\mu$ and $\sigma$ are outputs of the encoder. In this second approach, we feed the best design obtained so far throughout the optimization roll-out, initially starting with a uniformly distributed material. In such a way, the optimizer controls the perturbation of the best design so far, rather than the complete design.
In the results presented in the paper we use the larger architecture (VAE-L) with standard parameterization of the latent space (i.e controlling the full latent vector).

### v2

@@ E.1 Alternative configurations of the baseline VAE model
For the baseline VAE model, we tested 4 different configurations: 2 different architecture configurations (see Table 4), and 2 different parameterizations of the latent space. The results are summarized in Figure 3.
@@
Figure 26: Cumulative probability vs relative error (w.r.t. the MMA solution) for different optimization methods, evaluated on smaller testing sets (10 problems each, unlike on the full results), repeated with 10 different random seeds. The distribution does not add up to 1 due to cut-off at 300%.
@@
For optimization with VAE, we explore two different parameterizations of the latent space: 1) where the latent optimizer controls the entire input to the decoder ($\vec{z}$), and 2) where we utilize the information from the encoder, and the optimizer controls the additive term $\epsilon$, such that the input to the decoder $z=\mu+\sigma\epsilon$, where $\mu$ and $\sigma$ are outputs of the encoder. In this second approach, we feed the best design obtained so far throughout the optimization roll-out, initially starting with a uniformly distributed material. In such a way, the optimizer controls the perturbation of the best design so far, rather than the complete design. In the results presented in the article, we use the larger architecture (VAE-L) with standard parameterization of the latent space (i.e., controlling the full latent vector).
Appendix F Phase-field fracture simulation
Table 5: Detailed parameters of the phase-fields fracture simulation
Parameter value
Young’s modulus: E 210 GPa
Critical energy release rate: $G_{c}$ 2.7 kJ/$m^{2}$
Poisson ratio: $\nu$ 0.3
Phase field characteristic length scale: l 2 mm
Staggered scheme convergence tolerance 1e-5
Maximum number of
staggered scheme iterations
20
Newton solver tolerance 1e-6
Number of load increments 20
Number of elements along x 128
Number of elements along y 64